OpenStudy (megannicole51):

Does the series converge or diverge? n/(n^2+1) n=1 to infinity Studying for an exam I just need an answer on this one please. Professor decided theres no need for an answer key for our study sheet....idiot.

OpenStudy (anonymous):

Diverge. Since: \[ \frac{n}{n^2+1}\sim\frac{1}{n} \]And: \[ \sum_n\frac{1}{n}\to \infty \]

OpenStudy (megannicole51):

i used the limit comparison test and got 1 so it diverges because of the p-series that states p<or equal to 1 diverges. can i do it that way too?

OpenStudy (anonymous):

You can't quite use limit comparison since: \[ \frac{1}{n}>\frac{n}{n^2+1}=\frac{1}{n+\frac{1}{n}} \]

OpenStudy (megannicole51):

i think i did it wrong then haha i picked the wrong Bn thank you:)

OpenStudy (anonymous):

For this, you should probably make use of the ratio test. And, yup.

OpenStudy (megannicole51):

when can i NOT use the ratio test?

OpenStudy (anonymous):

Generally, some cases are fairly complicated, and, if you have a better way of doing it, then that's fine---also it's possible for the ratio test to be inconclusive, if, for example, you have: \[ \lim_{n\to\infty}\left|\frac{a_{n+1}}{a_n}\right|=1 \] Otherwise, you should probably make use of the ratio test. Unless, of course, you want to find some nice way of doing the problem, which there are always many.

OpenStudy (megannicole51):

awesome thank you:)

OpenStudy (anonymous):

Yup.